OpenStudy (khantahmina):

x^4-3x^2 +2 Write the equation of the line tangent to the graph at x=1

OpenStudy (khantahmina):

Also to put it out there I didnt learn derivatives sooo i cant use derivatives

OpenStudy (jiteshmeghwal9):

\[y=\frac{dy}{dx}_{(1,0)}x+c\]

OpenStudy (khantahmina):

just when i mentioned it @jiteshmeghwal9

OpenStudy (khantahmina):

how would u do this without derivatives and etc

OpenStudy (jiteshmeghwal9):

\[y=x_1^3x-3x_1x+2\]x_1=1

OpenStudy (khantahmina):

nope never learned that either

OpenStudy (jiteshmeghwal9):

maybe @FaiqRaees can help u

OpenStudy (khantahmina):

i learned about the diffferencce quotient formula

OpenStudy (khantahmina):

i learned it this way

OpenStudy (faiqraees):

What h and x represents in the answer

OpenStudy (khantahmina):

idk *sigh*

OpenStudy (khantahmina):

@FaiqRaees so basically no help i suppose

OpenStudy (khantahmina):

@jay

OpenStudy (jango_in_dtown):

a slight change @khantahmins... it will be lim h tends to 0 and then the expression you used..

OpenStudy (jango_in_dtown):

It represents the slope basically

OpenStudy (jango_in_dtown):

first let y=x^4-3x^2 +2

OpenStudy (jango_in_dtown):

then the slope at that point will be 4x^3-6x at x=1 i.e. -2

OpenStudy (jango_in_dtown):

the value of y at x=1 is 0

OpenStudy (jango_in_dtown):

so the required tangent line is y-0=-2(x-1) i.e. 2x+y=2

OpenStudy (jango_in_dtown):

@khantahmina

OpenStudy (khantahmina):

oh sorry @jango_IN_DTOWN i was a bit busy

OpenStudy (khantahmina):

i didnt get what u meant by "then the slope at that point will be 4x^3-6x at x=1 i.e. -2"

OpenStudy (khantahmina):

@zepdrix

zepdrix (zepdrix):

Hmm, I'm not really sure what else we can do besides taking the limit of the difference quotient. We have a fourth degree term, so it will be a tad bit of work expanding out the stuff. Shouldn't be too bad though.

zepdrix (zepdrix):

\[\rm \frac{d}{dx}(x^4-3x^2+2)=\lim_{h\to0}\frac{(x+h)^4-3(x+h)^2+2-(x^4-3x^2+2)}{h}\]

zepdrix (zepdrix):

Understand the first step, setting up the difference quotient?

OpenStudy (khantahmina):

yes i only understand that part

OpenStudy (khantahmina):

@zepdrix

OpenStudy (khantahmina):

so i expanded it all and this is what i got

OpenStudy (khantahmina):

\[[x^4+4x^3h+6x^@h^2+4xh^3+h^4-3(x^2+2xh+h^2)-2]-(x^4+3x^2+2)/h\]

zepdrix (zepdrix):

Mmm ya looks good. Distributing the -3 and the -1 respectively gives us this,\[\rm \lim_{h\to0}\frac{\color{orangered}{x^4+4x^3h+6x^2h^2+4xh^3+h^4-3x^2-6xh-6h^2-2}-x^4-3x^2-2}{h}\]

zepdrix (zepdrix):

So what's going to happen is, everything that doesn't contain an h should cancel out.

OpenStudy (khantahmina):

ok lemme work that out n ill tell u what i got

OpenStudy (khantahmina):

@zepdrix i think u got a mistake

OpenStudy (khantahmina):

it s supposed to be -3h^2

OpenStudy (khantahmina):

not -6h^2

zepdrix (zepdrix):

\[\rm \lim_{h\to0}\frac{\color{orangered}{x^4+4x^3h+6x^2h^2+4xh^3+h^4-3x^2-6xh-3h^2-2}-x^4-3x^2-2}{h}\]Ah good catch hehe

OpenStudy (khantahmina):

:)

OpenStudy (khantahmina):

x^4 cancels out 2 cancels out

zepdrix (zepdrix):

Oh we have a problem with the squared term. We forgot the negative in our original setup I think.\[[x^4+4x^3h+6x^@h^2+4xh^3+h^4-3(x^2+2xh+h^2)-2]-(x^4\color{red}{+}3x^2+2)/h\]

OpenStudy (khantahmina):

thats it

zepdrix (zepdrix):

\[\rm \lim_{h\to0}\frac{\color{orangered}{x^4+4x^3h+6x^2h^2+4xh^3+h^4-3x^2-6xh-3h^2-2}-x^4+3x^2-2}{h}\]So distributing the negative would give us +3x^2 like this.

OpenStudy (khantahmina):

noo the original had a -3x^2

zepdrix (zepdrix):

Yes, that's why I wrote it in red. Red is bad.

OpenStudy (khantahmina):

so distributing it would be +3x^2

OpenStudy (khantahmina):

got it

zepdrix (zepdrix):

So the squared terms cancel? Yay!

OpenStudy (khantahmina):

so \[4x63h+6x^2h+4xh^3+h^4-6xh-3h^2/h\]

OpenStudy (khantahmina):

4x^3h imeant ^^^

zepdrix (zepdrix):

Oh the carot is on the 6 key haha, I see what happened there ;o

zepdrix (zepdrix):

\[\large\rm \lim_{h\to0}\frac{4x^3h+6x^2h^2+4xh^3+h^4-6xh-3h^2}{h}\]Ok good. Next, divide an h out of everything, ya?

OpenStudy (khantahmina):

ok

zepdrix (zepdrix):

Squared h on your second term*

OpenStudy (khantahmina):

where

OpenStudy (khantahmina):

6x^2h^2?

zepdrix (zepdrix):

ya

OpenStudy (khantahmina):

so i have \[h(4x^3+6x^2+4xh^2+h^3-6x-3h) /h\]

OpenStudy (khantahmina):

ok imm a little confused here n impatient how is this going to help me find the equation of th e line tangent to x=1

zepdrix (zepdrix):

No no you forgot the h^2 on the second term again :O

zepdrix (zepdrix):

You should still have an h on that second term after factoring.\[h(4x^3+6x^2\color{royalblue}{h}+4xh^2+h^3-6x-3h) /h\]

OpenStudy (khantahmina):

im soo lost

zepdrix (zepdrix):

So let's talk about what the difference quotient is... It's just your slope formula. It let's you find the slope between some point x, and some other point which is distance h away from the first point.